An asset was purchased for $105,000 on January 1, Year 1 and originally estimated to have a useful life of 12 years with a residual value of $13,500. At the beginning of the third year, it was determined that the remaining useful life of the asset was only 4 years with a residual value of $2,000. Calculate the third-year depreciation expense using the revised amounts and straight-line method. Round your answer to the nearest dollar.
Answer #1

Answer--------Depreciation for year 3 = $21,938

Working

Straight line Method -For year 1 and 2
A Cost $ 105,000
B Residual Value $ 13,500
C=A - B Depreciable base $ 91,500
D Life [in years left ]                                12
E=C/D Annual SLM depreciation $ 7,625.00

.

Depreciation schedule-Straight line method- For first 2 complete years
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
1 $        105,000.00 $ 7,625.00 $ 7,625.00 $ 97,375.00
2 $ 97,375.00 $ 7,625.00 $ 15,250.00 $ 89,750.00

.

Straight line Method -For year 3 and onward
A Ending book value at Year 2 end $ 89,750
B Residual Value $ 2,000
C=A - B Depreciable base $ 87,750
D Life [in years left ]                                  4
E=C/D Annual SLM depreciation $ 21,938